OpenStudy (potatoe):

how to find the maclaurin series for the function f(x)=cosh(3x)

OpenStudy (anonymous):

\[ \large \cosh x= 1+ \frac{x^2}{2!} +\frac{x^4}{4!}+\frac{x^6}{6!} \cdots \]

OpenStudy (anonymous):

\[ \large \cosh x= \sum_{n=1}^\infty \frac{x^{2n}}{(2n)!} \] General case, you can use substitution from here.

OpenStudy (mathmate):

Cosh(x)=(e^x+e^(-x))/2 Starting from e^x=1+x+x^2/2!+x^3/3!+... You would add the two series term by term, and will find that the odd powers cancel out to get the series given by Spacelimbus above.

OpenStudy (anonymous):

Oh, I believe I misunderstood this question maybe, if you need to derive it from the beginning, then you clearly want to use the way that @mathmate suggest above, using the definition of \(\cosh x\).

OpenStudy (potatoe):

it wouldn't affect the problem being cosh3x rather than cosh x?

OpenStudy (potatoe):

never mnd i got it

OpenStudy (potatoe):

thank you very much though

OpenStudy (potatoe):

calc 2 = death of me

OpenStudy (mathmate):

You'd replace x in the series by 3x, which means that you have cosh(3x)=1+(3x)^2/2! + (3x)^4/4! + ... which reduces to: cosh(3x)=1+3^2 x^2/2! + 3^4 x^4/4! + ...

OpenStudy (mathmate):

Don't worry, working out a few problems like this will do magic for you exams.